Carnivores may oxidize amino acids as a source of energy. Terrestrial vertebrates and sharks oxidize amino...

  1. Carnivores may oxidize amino acids as a source of energy. Terrestrial vertebrates and sharks oxidize amino acids to urea (NH2CONH2), carbon dioxide and water - as an oxidation product, urea has a low toxicity, high solubility and easy to eliminate from the body. An example is provided by glycine, NH2CH2CO2H. At 37 °C and 1.00 atm, what volume of carbon dioxide is produced for each microgram of glycine produced? Many aquatic venerates oxidize amino acids to ammonia, carbon dioxide and water. For each microgram of glycine oxidized, what volume of carbon dioxide is produced at 25 °C and 1.00 atm.
Answer #1

Write the balanced chemical equation as

2 NH2CH2CO2H + 3 O2 -----------> NH2CONH2 + 3 CO2 + 3 H2O

As per the stoichiometric equation,

2 moles glycine = 3 moles CO2.

The atomic masses are

H: 1.008 g/mol.

C: 12.011 g/mol.

N: 14.007 g/mol.

O: 15.999 g/mol.

Molar mass of glycine, NH2CH2CO2H = (1*14.007 + 5*1.008 + 2*12.011 + 2*15.999) g/mol

= 75.067 g/mol.

Mass of glycine = 1.0 µg

= (1.0 µg)*(1 g)/(1.0*106 µg)

= 1.0*10-6 g.

Mol(s) glycine corresponding to 1.0 µg glycine = (1.0*10-6 g)/(75.067 g/mol)

= 1.332*10-8 mol.

Mol(s) CO2 corresponding to 1.0 µg glycine = (1.332*10-8 mol glycine)*(3 mols CO2)/(2 mols glycine)

= 1.998*10-8 mol.

Assume the liberated CO2 to behave as an ideal gas.

Pressure of the gas = 1.00 atm.

Temperature of the gas = 37ºC = (37 + 273) K = 310 K.

Use the ideal gas law to determine the volume of CO2 liberated.

PV = nRT

where n denotes the number of moles of the gas.

Therefore,

V = nRT/P

Plug in values and get

V = (1.998*10-8 mol)*(0.082 L-atm/mol.K)*(310 K)/(1.00 atm)

= 5.0789*10-7 L

= (5.0789*10-7 L)*(1.0*106 µL)/(1 L)

= 0.50789 µL

≈ 0.51 µL.

The volume of CO2 gas evolved per microgram of glycine oxidized = 0.51 µL (ans).